OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. A. y = -2x + 14 B. y = -2x + 5 C. y = -2x + 9 D. y = -2x + 6

jimthompson5910 (jim_thompson5910):

y - 4 = -2(x - 5) y - 4 = -2x + 10 y = -2x + 10 + 4 y = -2x + 14